OpenStudy (lgbasallote):

\[\int \int_R \sin (x - y) dA. R = [(x,y) | 0 \le x \le \pi / 2, 0 \le y \le \pi / 2]\] can someone tell me what that meant o.O

OpenStudy (anonymous):

its a double integral

OpenStudy (anonymous):

the limits for x and y are provided, it straight forward

OpenStudy (anonymous):

this is mulitvariable calculus, its not something that would show up in calc 2 or integral calc

OpenStudy (lgbasallote):

what does "multivariable calculus" mean exactly?

OpenStudy (lgbasallote):

and why is it not under integral calculus?

OpenStudy (zarkon):

latex tip...use \[\iint\] \iint when writing a double integral

OpenStudy (lgbasallote):

ohh...was it that obvious haha

OpenStudy (zarkon):

\[\iint_R \sin (x - y) dA\,\,\,\,\,\,\,\,\,\,\, R = \{(x,y) | 0 \le x \le \pi / 2, 0 \le y \le \pi / 2\}\]

OpenStudy (zarkon):

\[=\int\limits_{0}^{\pi/2} \int\limits_{0}^{\pi/2} \sin (x - y) dx\,dy\]

OpenStudy (lgbasallote):

but what does R mean? i mean wht it symbolizes

OpenStudy (zarkon):

\(R\) is the region is the \(xy\) plane that you are integrating over.

OpenStudy (lgbasallote):

and dA? differential area?

OpenStudy (zarkon):

yes...typically it is written \(dA=dx dy=dydx\)

OpenStudy (lgbasallote):

ohhh thanks :DD